i have 16 flowers plan on using 5 differnt ones. how many different selections of 5 can i make?
Answer by stanbon(75887)

i have 16 flowers plan on using 5 differnt ones. how many different selections of 5 can i make?
----
Ans: 16C5 = 16!/(16-5)!*5! = (16*15*14*13*12)/(1*2*3*4*5) = 4368
-----------------
Cheers,
Stan H.
===================